# Large Language Models for Multilingual Vulnerability Detection: How Far Are We?

### v1

@@ 3.2. Experimented Language Models
  • CodeT5P (Wang et al., 2023a) extends CodeT5’s capabilities through an architecture featuring a shallow encoder paired with a deep decoder. The training process follows a multi-stage approach, beginning with unimodal training before progressing to bimodal data integration.
@@
We focused exclusively on PLMs rather than employing GNNs or RNNs for multilingual vulnerability detection. This choice was motivated by PLMs’ proven strengths in vulnerability detection compared to RNNs and GNNs (Steenhoek et al., 2023). Moreover, alternative approaches (Zhang et al., 2023b; Liu et al., 2024; Steenhoek et al., 2024) heavily rely on Abstract Syntax Trees (ASTs). Since ASTs are inherently language-specific structures, creating a unified AST representation across different programming languages is impractical due to varying syntax and semantic constructs.
For LLMs, we studied five state-of-the-art models, including three open-source LLMs and two closed-source LLMs, which have demonstrated promising performance in various code-related tasks (Zhou et al., 2024a; Hou et al., 2023). The details of these LLMs are as follows:
@@
Similarly, we excluded LLM-based approaches like MSIVD (Yang et al., 2024) and GRACE (Lu et al., 2024) because they require AST construction, which is impractical for implementing unified ASTs across multiple programming languages and does not support line-level detection. We also omitted Vul-RAG (Du et al., 2024) because it cannot support or be extended to line-level vulnerability detection. As a result, we focused primarily on vanilla general and code-specific LLMs for multilingual vulnerability detection, following previous empirical studies (Fu et al., 2023a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c).
@@
((a))
((b))
Figure 2. Zero-shot prompt template for multilingual vulnerability detection
3.3. Strategies for Large Language Models
@@
  • Zero-shot prompting strategy: We developed a prompt structure with system and user roles, following prior works (Zhou et al., 2024b; Zhang et al., 2023a). We directly employed a structured instruction and a vulnerable function to prompt LLMs for multilingual vulnerability detection at both function and line levels, without using examples. Figure 2 shows the zero-shot prompt templates for function-level and line-level multilingual vulnerability detection. For function-level detection, we specify that the input is a code function written in a particular programming language, then ask the LLM to determine whether the function is vulnerable or clean. For line-level detection, we specify that the input is a vulnerable function written in a particular programming language, then ask the LLM to identify the vulnerable line number and its corresponding code.
    ((a))
    ((b))
    Figure 3. Few-shot prompt template for multilingual vulnerability detection
@@
    For function-level detection, an instruction-filled function pair contains three elements: a zero-shot prompt instruction (see Figure 2(a)), a given function, and a label indicating whether that function is vulnerable or clean. Using all 16,126 instances from the function-level training set, we constructed a function-level instruction-filled fine-tuning set. We then fine-tuned the LLMs on this fine-tuning set to detect function-level multilingual vulnerability using either zero-shot or few-shot prompts. Similarly, for line-level detection, an instruction-filled function pair consists of a zero-shot prompt instruction (see Figure 2(b)), a vulnerable function, and corresponding line labels. Using all 6,513 instances from the line-level training set, we created a line-level instruction-filled fine-tuning set. Then, we fine-tuned the LLMs on this set to detect line-level multilingual vulnerability using either zero-shot or few-shot prompts.
3.4. Metrics
@@
Following existing work (Fu and Tantithamthavorn, 2022; Steenhoek et al., 2024), we adopted four widely-used metrics, Accuracy, Precision, Recall, and F1-score, to assess all our studied models’ effectiveness in detecting function-level multilingual vulnerability as a binary classification task. Accuracy measures the proportion of correctly predicted instances out of all instances. Precision quantifies how many of the predicted positive instances are actually correct, while Recall assesses the proportion of actual positive instances correctly predicted by the model. The F1-score combines precision and recall into a single metric, providing a balanced view especially useful when dealing with imbalanced datasets. Specifically, in our study, we used binary-averaged Precision, Recall, and F1-score because function-level vulnerability detection can be formulated as a binary classification task. Since the ratio between positive (vulnerable) and negative (non-vulnerable) instances in the function-level dataset is approximately 50:50 (see Table 1), we considered Accuracy a critical metric for measuring the models’ effectiveness in multilingual vulnerability detection.
For line-level multilingual vulnerability detection, we formulated the line-level detection task as a binary classification problem for each code line by following the existing work (Li et al., 2021b). Therefore, we also adopted Accuracy, Precision, Recall, and F1-score to assess all our studied models’ effectiveness in detecting line-level multilingual vulnerability.
@@
Furthermore, we adopted the False Positive Rate (FPR) to measure how often a classifier incorrectly predicts the positive class when the actual class is negative, and False Negative Rate (FNR) to quantify how often the model mistakenly classifies positive instances as negative. In vulnerability detection, false positives (when clean code is flagged as vulnerable) cause developers to waste resources investigating secure code, while false negatives (when vulnerable code is flagged as clean) pose a fatal risk by leaving security flaws unaddressed. The trade-off between FPR and FNR, achieving lower FNR at the cost of lower FPR, is critical for measuring the intrinsic effectiveness of PLMs and LLMs in multilingual vulnerability detection. Specifically, FPR is defined as $\frac{FP}{FP+TN}$ where FP represents false positives and TN represents true negatives. FNR is defined as $\frac{FN}{FN+TP}$ where FN represents false negatives and TP represents true positives. A high FPR means many clean functions and lines remain unrecognized. A high FNR means many vulnerable functions or lines remain undetected.
3.5. Implementation and Environment
@@
For pre-trained language models (PLMs), we replicated CodeBert, CodeT5, CodeT5P, UniXCoder, LineVul, and Text-Embedding-Models following existing works (Steenhoek et al., 2024). We obtained all open-source pre-trained checkpoints (i.e., CodeBert, CodeT5, CodeT5P, UniXCoder, and LineVul) from Huggingface (Wolf et al., 2019). Since Text-Embedding-Models are a series of closed-source PLMs, we accessed them through OpenAI’s API. For our experiments with closed-source LLMs, we accessed GPT-3.5-Turbo (model version gpt-3.5-turbo-0125) and GPT-4o (model version gpt-4o-2024-08-06) through OpenAI’s APIs (OpenAI, 2024a). For open-source LLMs, we utilized pre-trained checkpoints from Huggingface (Wolf et al., 2019) for DeepSeek-Coder (6.7B parameters), Code Llama (7B parameters), and Llama 3 (8B parameters). We implemented Low-Rank Adaptation (LoRA (Hu et al., 2021)) during fine-tuning to optimize computational efficiency and prevent overfitting. We conducted all the experiments on an Intel Xeon CPU Gold-6342 machine with 512 GB RAM, Ubuntu 20.04.6, and two A800 GPUs. More implementation details can be found in our replication package. To facilitate future research, we have made all the used datasets and model execution scripts publicly available.
4. Evaluation Results
@@
4.1. RQ1: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the function-level?
@@
Regarding PLMs, they are pre-trained on a corpus of natural language and source code snippets from various programming languages, but they have not been exposed to the task of function-level vulnerability detection, particularly across multilingual vulnerability. Prior studies (Fu and Tantithamthavorn, 2022; Steenhoek et al., 2023, 2024) have demonstrated the effectiveness of fine-tuned PLMs for function-level vulnerability detection. Therefore, we fine-tune all studied open-source PLMs (i.e., CodeBert, CodeT5, CodeT5P, UniXCoder, and LineVul) on the function-level training set. The model architecture for function-level multilingual vulnerability detection combines a PLM’s encoder with a downstream dense layer that acts as a binary classifier. We train all parameters of both the encoder and binary classifier through supervised learning on the function-level training set. The trained encoder outputs a code embedding representing the given input given code. This embedding is then fed into the trained binary classifier, which produces a binomial distribution indicating the probability of the given function being vulnerable. For closed-source PLMs (i.e., Text-Embedding-Ada-002, Text-Embedding-3-Small, and Text-Embedding-3-Large), we follow the same training and inference approach. Since these PLMs do not allow access to their parameters, we treat them as frozen encoders and only train the binary classifier.
@@
To evaluate the effectiveness of the studied PLMs and LLMs in detecting function-level multilingual vulnerabilities, we employed multiple metrics, including Accuracy, Precision, Recall, F1-score, FPR, and FNR as detailed in Section 3.4. Given the balanced distribution between positive and negative labels, Accuracy was used as the primary evaluation metric.
@@
Table 3. Performance of PLMs and LLMs for function-level multilingual vulnerability detection
@@
Techniques Accuracy Recall Precision F1-score FPR FNR Dummy Classifiers ${DummyClf}_{vul}$ 0.5030 1.0000 0.5030 0.6693 1.0000 0.0000 ${DummyClf}_{clean}$ 0.4970 0.0000 0.0000 0.0000 0.0000 1.0000 PLMs Text-Embedding-3-Large 0.5064 0.9696 0.5049 0.6640 0.9623 0.0304 Text-Embedding-3-Small 0.5044 0.9676 0.5038 0.6626 0.9643 0.0324 Text-Embedding-Ada-002 0.5030 1.0000 0.5030 0.6693 1.0000 0.0000 CodeBERT 0.5173 1.0000 0.5103 0.6757 0.9712 0.0000 LineVul 0.5173 1.0000 0.5103 0.6757 0.9712 0.0000 UniXcoder 0.5814 0.8930 0.5518 0.6822 0.7339 0.1070 CodeT5 0.5854 0.9342 0.5519 0.6939 0.7676 0.0658 CodeT5P 0.6037 0.9529 0.5626 0.7075 0.7496 0.0471 LLMs (Zero-Shot Prompting) DeepSeek-Coder 0.5005 0.1796 0.5097 0.2656 0.1748 0.8204 Code Llama 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 Llama 3 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 GPT-3.5-Turbo 0.4827 0.6183 0.4888 0.5459 0.6544 0.3817 GPT-4o 0.4985 0.3631 0.5020 0.4214 0.2632 0.7399 LLMs (Few-Shot Prompting) DeepSeek-Coder 0.3870 0.5309 0.4146 0.4656 0.7587 0.4691 Code Llama 0.4763 0.6251 0.4840 0.5456 0.6743 0.3749 Llama 3 0.4299 0.5260 0.4437 0.4814 0.6673 0.4740 GPT-3.5-Turbo 0.3618 0.4426 0.3835 0.4109 0.7200 0.5574 GPT-4o 0.5035 0.4347 0.5074 0.4683 0.4270 0.5653 LLMs (Instruction Tuning + Zero-Shot Prompting) DeepSeek-Coder 0.5039 0.3023 0.5116 0.3800 0.2920 0.6977 Code Llama 0.4783 0.5103 0.4824 0.4959 0.5531 0.4897 Llama 3 0.5192 0.5348 0.5215 0.5281 0.4965 0.4652 GPT-3.5-Turbo 0.5454 0.5132 0.5517 0.5318 0.4220 0.4868 GPT-4o 0.5874 0.4622 0.6206 0.5298 0.2859 0.5378 LLMs (Instruction Tuning + Few-Shot Prompting) DeepSeek-Coder 0.4906 0.4789 0.4934 0.4861 0.4975 0.5201 Code Llama 0.5094 0.6408 0.5098 0.5678 0.6226 0.3592 Llama 3 0.4946 0.4848 0.4975 0.4911 0.4955 0.5152 GPT-3.5-Turbo 0.5000 0.4524 0.5033 0.4765 0.4518 0.5475 GPT-4o 0.7196 0.6722 0.7454 0.7069 0.2324 0.3278
@@
Results. Table 3 shows the comparison results among PLMs and LLMs in terms of average Accuracy, Recall, Precision, F1-score, FPR, and FNR in detecting function-level multilingual vulnerability. Within each metric, the values in bold indicate the model that exhibits the best performance among all PLMs and LLMs. Figure 4 presents the performance of PLMs and LLMs with instruction tuning and few-shot prompting in seven programming languages in terms of Accuracy.
@@
Observation 1: Among PLMs, CodeT5P exhibits superior performance in function-level multilingual vulnerability detection. In the PLMs’ category, text-embedding models like Text-Embedding-3-Large, Text-Embedding-3-Small, and Text-Embedding-Ada-002 all hover around 0.5000 in Accuracy, indicating near-random performance on a balanced dataset. Despite high Recall scores (all >0.9600), their low Precision (around 0.5000) results in modest F1-scores (around 0.6600). This suggests a bias toward over-identifying functions as vulnerable, in other words, a high FPR (all >0.9600). Similarly, CodeBERT and LineVul achieve slightly higher Accuracy (0.5163 and 0.5173) and perfect Recall (1.0000) but suffer from low Precision (0.5098 and 0.5103). This leads to middling F1-scores (0.6753 and 0.6757) and limited practical utility, as they cannot reliably discriminate between vulnerable and non-vulnerable functions.
@@
More modern code-specific models show improved balance. UniXcoder and CodeT5 both outperform general text-embedding models and simple code-specific models, with UniXcoder reaching 0.5814 Accuracy and CodeT5 achieving 0.5854. These models strike a better trade-off between Recall and Precision, resulting in higher F1-scores (0.6822 and 0.6939, respectively). Most notably, CodeT5P delivers the strongest performance across all PLMs, achieving the highest Accuracy of 0.6037 along with a solid F1-score of 0.7075. Furthermore, the CodeT5P reaches FPR of 0.7496 and FNR of 0.0471, which implies that although CodeT5P can detect more vulnerable functions, it comes at the cost of wrongly detecting more clean functions. This indicates that CodeT5P, which incorporates identifier-aware training objectives and enriched representation learning, is particularly effective at capturing the multilingual semantic nuances required for function-level multilingual vulnerability detection. Overall, the results demonstrate that while traditional PLMs struggle with precision and overall Accuracy, task-aware models like CodeT5P offer significant improvements, making them more suitable for multilingual vulnerability detection scenarios at the function-level.

### v2

@@ 2.2 Experimented Language Models
  • CodeT5P (Wang et al., 2023b) extends CodeT5’s capabilities through an architecture featuring a shallow encoder paired with a deep decoder. The training process follows a multi-stage approach, beginning with unimodal training before progressing to bimodal data integration.
@@
We focused exclusively on PLMs rather than employing GNNs or RNNs for multilingual vulnerability detection. This choice was motivated by PLMs’ proven strengths in vulnerability detection compared to RNNs and GNNs (Steenhoek et al., 2023). Moreover, alternative approaches (Zhang et al., 2023a; Liu et al., 2024; Steenhoek et al., 2024a) heavily rely on Abstract Syntax Trees (ASTs). Since ASTs are inherently language-specific structures, creating a unified AST representation across different programming languages is impractical due to varying syntax and semantic constructs.
For LLMs, we studied five state-of-the-art models, including three open-source LLMs and two closed-source LLMs, which have demonstrated promising performance in various code-related tasks (Zhou et al., 2024a; Hou et al., 2023). We strategically selected three open-source LLMs with parameter counts ranging from 6.7B to 8B. This choice is driven by pragmatic considerations: many practitioners lack the high-end computational resources required to deploy or instruction-tune larger models. Furthermore, our selection is supported by recent findings (Alizadeh et al., 2025; Zhong et al., 2025) indicating that increased model scale does not consistently guarantee superior performance on specialized tasks. Consequently, these models represent an optimal balance between computational efficiency and competitive predictive capability. The details of these LLMs are as follows:
@@
To ensure a rigorous comparison against the state-of-the-art, we extend our evaluation to include GRACE (Lu et al., 2024), a representative framework that enhances LLMs by injecting structural program information derived from ASTs. Given that GRACE utilizes a zero-shot prompting strategy to elicit vulnerability insights without task-specific parameter updates, we categorize it as a zero-shot prompting baseline in our study. We exclude MSIVD (Du et al., 2024) because we lack attack complexity data and precise vulnerability explanations across multiple languages, which could lead to inconsistent instructional quality across languages. Similarly, we exclude Vul-RAG (Yang et al., 2024a) to avoid knowledge availability bias, as we lack sufficient language-specific information to construct the vulnerability knowledge base. As a result, we focused primarily on vanilla general and code-specific LLMs for multilingual vulnerability detection, following previous empirical studies (Fu et al., 2023a; Yin et al., 2024; Yin, 2024; Zhou et al., 2024c).
@@
Figure 2: Zero-shot prompt template for multilingual vulnerability detection
2.3 Strategies for Large Language Models
@@
  • Zero-shot prompting strategy: We developed a prompt structure with system and user roles, following prior works (Zhou et al., 2024b; Zhang et al., 2023b). We directly employed a structured instruction and a vulnerable function to prompt LLMs for multilingual vulnerability detection at both function and line levels, without using examples. Figure 2 shows the zero-shot prompt templates for function-level and line-level multilingual vulnerability detection. For function-level detection, we specify that the input is a code function written in a particular programming language, then ask the LLM to determine whether the function is vulnerable or clean. For line-level detection, we specify that the input is a vulnerable function written in a particular programming language, then ask the LLM to identify the vulnerable line number and its corresponding code.
    Figure 3: Few-shot prompt template for multilingual vulnerability detection
@@
    For function-level detection, an instruction-filled function pair contains three elements: a zero-shot prompt instruction (see Figure 2(a)), a given function, and a label indicating whether that function is vulnerable or clean. Using all 16,126 instances from the function-level training set, we constructed a function-level instruction-filled fine-tuning set. We then fine-tuned the LLMs on this fine-tuning set to detect function-level multilingual vulnerability using either zero-shot or few-shot prompts. Similarly, for line-level detection, an instruction-filled function pair consists of a zero-shot prompt instruction (see Figure 2(b)), a vulnerable function, and corresponding line labels. Using all 6,513 instances from the line-level training set, we created a line-level instruction-filled fine-tuning set. Then, we fine-tuned the LLMs to detect line-level multilingual vulnerability using either zero-shot or few-shot prompts.
    In our study, we adopt Parameter-Efficient Fine-Tuning (PEFT) via Low-Rank Adaptation (LoRA) (Hu et al., 2021) to perform instruction tuning on LLMs on both function- and line-level multilingual vulnerability detection. LoRA reduces the number of trainable parameters by freezing the original model weights and injecting trainable low-rank matrices into each target weight matrix. Specifically, given an original weight matrix $W_{0}\in\mathbb{R}^{d\times k}$, LoRA represents the weight update $\Delta W$ as the product of two low-rank matrices $A\in\mathbb{R}^{d\times r}$ and $B\in\mathbb{R}^{r\times k}$, i.e.,
    $\Delta W=AB,\quad r\ll\min(d,k),$ (1)
    and the adapted weight is
    $W=W_{0}+\alpha\cdot\Delta W,$ (2)
    where $\alpha$ is a scaling factor. This formulation enables us to fine-tune large models for instruction-following tasks with significantly reduced computational cost while maintaining competitive performance. Moreover, its parameter efficiency allows us to adapt the same base model across multiple programming languages and domains without requiring full retraining, which is particularly advantageous for multilingual and cross-domain instruction tuning.
2.4 Metrics
@@
Following existing work (Fu and Tantithamthavorn, 2022; Steenhoek et al., 2024a), we adopted four widely-used metrics, Accuracy, Precision, Recall, and F1-score, to assess all our studied models’ effectiveness in detecting function-level multilingual vulnerability as a binary classification task. Accuracy measures the proportion of correctly predicted instances out of all instances. Precision quantifies how many of the predicted positive instances are actually correct, while Recall assesses the proportion of actual positive instances correctly predicted by the model. The F1-score combines precision and recall into a single metric, providing a balanced view especially useful when dealing with imbalanced datasets. Specifically, in our study, we used binary-averaged Precision, Recall, and F1-score because function-level vulnerability detection can be formulated as a binary classification task. For line-level multilingual vulnerability detection, we formulated the line-level detection task as a binary classification problem for each code line by following the existing work (Ding et al., 2024b). Therefore, we also adopted Accuracy, Precision, Recall, and F1-score to assess all our studied models’ effectiveness in detecting line-level multilingual vulnerability.
@@
Furthermore, we adopted the False Positive Rate (FPR) to measure how often a classifier incorrectly predicts the positive class when the actual class is negative, and False Negative Rate (FNR) to quantify how often the model mistakenly classifies positive instances as negative. In vulnerability detection, false positives (when clean code is flagged as vulnerable) cause developers to waste resources investigating secure code, while false negatives (when vulnerable code is flagged as clean) pose a fatal risk by leaving security flaws unaddressed. The trade-off between FPR and FNR, achieving lower FNR at the cost of lower FPR, is critical for measuring the intrinsic effectiveness of PLMs and LLMs. Specifically, FPR is defined as $\frac{FP}{FP+TN}$ where FP represents false positives and TN represents true negatives. FNR is defined as $\frac{FN}{FN+TP}$ where FN represents false negatives and TP represents true positives. A high FPR means many clean functions and lines remain unrecognized. A high FNR means many vulnerable functions or lines remain undetected.
Following the recommendations of Uddin et al. (2025), we adopt the Matthews Correlation Coefficient (MCC) and Area Under the Receiver Operating Characteristic Curve (AUC) to provide a comprehensive evaluation of multilingual vulnerability detection at both the function and line levels. MCC serves as a robust measure of binary classification quality by accounting for all four quadrants of the confusion matrix, calculated as $\frac{TP\times TN-FP\times FN}{\sqrt{(TP+FP)(TP+FN)(TN+FP)(TN+FN)}}$. Complementing this, AUC is employed to evaluate the model’s discriminative ability across all possible classification thresholds. Statistically, AUC represents the probability that a randomly selected vulnerable instance will be ranked higher than a benign one, making it invariant to the specific choice of threshold. We compute the AUC by using the trapezoidal rule.
To reduce ambiguity in calculating metrics across multiple programming languages, we use Accuracy as an illustrative example. Suppose our test set consists of only two languages: Python and Java. Java contains one sample and Python contains two samples. A model correctly predicts one Java sample and one Python sample. Instead of calculating Accuracy for each language separately (Python: 50%, Java: 100%) and then averaging those percentages (75%), we aggregate all samples first. The Total Correct Predictions is two and the Total Samples is three. The global Accuracy is calculated as $\text{Accuracy}=\frac{\text{Total Correct Predictions}}{\text{Total Samples}}=\frac{2}{3}\approx 66.7\%$. This approach ensures that performance is weighted by the actual distribution of samples in each language, providing an objective measurement of the holistic performance of PLMs and LLMs.
2.5 Implementation and Environment
@@
For pre-trained language models (PLMs), we replicated CodeBert, CodeT5, CodeT5P, UniXCoder, LineVul, and Text-Embedding-Models following existing works (Steenhoek et al., 2024a). We obtained all open-source pre-trained checkpoints (i.e., CodeBert, CodeT5, CodeT5P, UniXCoder, and LineVul) from Huggingface (Wolf et al., 2019). Since Text-Embedding-Models are a series of closed-source PLMs, we accessed them through OpenAI’s API. For our experiments with closed-source LLMs, we accessed GPT-3.5-Turbo (model version gpt-3.5-turbo-0125) and GPT-4o (model version gpt-4o-2024-08-06) through OpenAI’s APIs (OpenAI, 2024a). For open-source LLMs, we utilized pre-trained checkpoints from Huggingface (Wolf et al., 2019) for DeepSeek-Coder (6.7B parameters), Code Llama (7B parameters), and Llama 3 (8B parameters). We implemented Low-Rank Adaptation during fine-tuning to optimize computational efficiency and prevent overfitting. We conducted all the experiments on an Intel Xeon CPU Gold-6342 machine with 512 GB RAM, Ubuntu 20.04.6, and two A800 GPUs. More implementation details can be found in our replication package. To facilitate future research, we have made all the used datasets and model execution scripts publicly available.
3 Evaluation Results
@@
3.1 RQ1: How effective are PLMs and LLMs in detecting multilingual vulnerabilities at the function-level?
@@
Regarding PLMs, they are pre-trained on a corpus of natural language and source code snippets from various programming languages, but they have not been exposed to the task of function-level vulnerability detection, particularly across multilingual vulnerability. Prior studies (Fu and Tantithamthavorn, 2022; Steenhoek et al., 2023, 2024a) have demonstrated the effectiveness of fine-tuned PLMs for function-level vulnerability detection. Therefore, we fine-tune all studied open-source PLMs (i.e., CodeBert, CodeT5, CodeT5P, UniXCoder, and LineVul) on the function-level training set. The model architecture for function-level multilingual vulnerability detection employs a hybrid design that integrates a PLM encoder with a specialized binary classification head. Specifically, the binary classifier is implemented as a multi-layer perceptron (MLP) that operates on the aggregate sequence representation. To facilitate non-linear mapping and mitigate overfitting, the architecture incorporates a hyperbolic tangent ($\tanh$) activation function interleaved with stochastic dropout layers. The forward pass begins by extracting the representation of the leading sequence token (equivalent to the [CLS] or <s> token), which is then processed through a dense layer and non-linearities. The final output of this classification head is the raw logit, representing the model’s unnormalized confidence score before being mapped to the binary prediction space. We train all parameters of both the encoder and binary classifier through supervised learning on the function-level training set. The trained encoder outputs a code embedding representing the given input given code. This embedding is then fed into the trained binary classifier, which produces a binomial distribution indicating the probability of the given function being vulnerable. For closed-source PLMs (i.e., Text-Embedding-Ada-002, Text-Embedding-3-Small, and Text-Embedding-3-Large), we follow the same training and inference approach. Since these PLMs do not allow access to their parameters, we treat them as frozen encoders and only train the binary classifier.
@@
To evaluate the effectiveness of the studied PLMs and LLMs in detecting function-level multilingual vulnerabilities, we employed multiple metrics, including Accuracy, Precision, Recall, F1-score, FPR, FNR, MCC, and AUC as detailed in Section 2.4. Given the balanced distribution between positive and negative labels, Accuracy was used as the primary evaluation metric.
@@
Table 3: Performance of PLMs and LLMs for function-level multilingual vulnerability detection on balance scenario
@@
Techniques Accuracy Recall Precision F1-score FPR FNR MCC AUC Dummy Classifiers ${DummyClf}_{vul}$ 0.5030 1.0000 0.5030 0.6693 1.0000 0.0000 0.0000 0.5000 ${DummyClf}_{clean}$ 0.4970 0.0000 0.0000 0.0000 0.0000 1.0000 0.0000 0.5000 PLMs Text-Embedding-3-Large 0.5064 0.9696 0.5049 0.6640 0.9623 0.0304 0.0201 0.5037 Text-Embedding-3-Small 0.5044 0.9676 0.5038 0.6626 0.9643 0.0324 0.0093 0.5016 Text-Embedding-Ada-002 0.5030 1.0000 0.5030 0.6693 1.0000 0.0000 0.0000 0.5000 CodeBERT 0.5173 1.0000 0.5103 0.6757 0.9712 0.0000 0.1212 0.5144 LineVul 0.5173 1.0000 0.5103 0.6757 0.9712 0.0000 0.1212 0.5144 UniXcoder 0.5814 0.8930 0.5518 0.6822 0.7339 0.1070 0.2045 0.5796 CodeT5 0.5854 0.9342 0.5519 0.6939 0.7676 0.0658 0.2342 0.5833 CodeT5P 0.6037 0.9529 0.5626 0.7075 0.7496 0.0471 0.2860 0.6015 LLMs (Zero-Shot Prompting) DeepSeek-Coder 0.5005 0.1796 0.5097 0.2656 0.1748 0.8204 0.0063 0.5024 Code Llama 0.4877 0.9156 0.4950 0.6426 0.9454 0.0844 -0.0585 0.4851 Llama 3 0.5005 0.5015 0.5034 0.5025 0.5005 0.4985 0.0010 0.5005 GPT-3.5-Turbo 0.4827 0.6183 0.4888 0.5459 0.6544 0.3817 -0.0375 0.4819 GPT-4o 0.4985 0.3631 0.5020 0.4214 0.2632 0.7399 -0.0035 0.4985 GRACE 0.5061 0.1673 0.5323 0.2546 0.1495 0.8327 0.0244 0.5089 LLMs (Few-Shot Prompting) DeepSeek-Coder 0.3870 0.5309 0.4146 0.4656 0.7587 0.4691 -0.2378 0.3861 Code Llama 0.4763 0.6251 0.4840 0.5456 0.6743 0.3749 -0.0515 0.4754 Llama 3 0.4299 0.5260 0.4437 0.4814 0.6673 0.4740 -0.1440 0.4293 GPT-3.5-Turbo 0.3618 0.4426 0.3835 0.4109 0.7200 0.5574 -0.2810 0.3613 GPT-4o 0.5035 0.4347 0.5074 0.4683 0.4270 0.5653 0.0078 0.5039 LLMs (Instruction Tuning + Zero-Shot Prompting) DeepSeek-Coder 0.5039 0.3023 0.5116 0.3800 0.2920 0.6977 0.0113 0.5051 Code Llama 0.4783 0.5103 0.4824 0.4959 0.5531 0.4897 -0.0429 0.5091 Llama 3 0.5192 0.5348 0.5215 0.5281 0.4965 0.4652 0.0383 0.5192 GPT-3.5-Turbo 0.5454 0.5132 0.5517 0.5318 0.4220 0.4868 0.0914 0.5456 GPT-4o 0.5874 0.4622 0.6206 0.5298 0.2859 0.5378 0.1820 0.5881 LLMs (Instruction Tuning + Few-Shot Prompting) DeepSeek-Coder 0.4906 0.4789 0.4934 0.4861 0.4975 0.5201 -0.0176 0.4912 Code Llama 0.5094 0.6408 0.5098 0.5678 0.6226 0.3592 0.0188 0.5091 Llama 3 0.4946 0.4848 0.4975 0.4911 0.4955 0.5152 -0.0107 0.4946 GPT-3.5-Turbo 0.5000 0.4524 0.5033 0.4765 0.4518 0.5475 0.0005 0.5003 GPT-4o 0.7196 0.6722 0.7454 0.7069 0.2324 0.3278 0.4360 0.7169
@@
Results. Table 3 shows the comparison results among PLMs and LLMs in terms of average Accuracy, Recall, Precision, F1-score, FPR, FNR, MCC, and AUC in detecting function-level multilingual vulnerability. Within each metric, the values in bold indicate the model that exhibits the best performance among all PLMs and LLMs. Figure 4 presents the performance of PLMs and LLMs with instruction tuning and few-shot prompting in seven programming languages in terms of Accuracy.
@@
Observation 1: Among PLMs, CodeT5P exhibits superior performance in function-level multilingual vulnerability detection. In the PLMs’ category, text-embedding models like Text-Embedding-3-Large, Text-Embedding-3-Small, and Text-Embedding-Ada-002 all hover around 0.5000 in Accuracy, indicating near-random performance on a balanced dataset. Despite high Recall scores (all >0.9600), their low Precision (around 0.5000) results in modest F1-scores (around 0.6600). This suggests a bias toward over-identifying functions as vulnerable, in other words, a high FPR (all >0.9600). The MCC scores for these models are consistently near zero (e.g., 0.0201 for Text-Embedding-3-Large), and AUC values hover around 0.5 (e.g., 0.5037), further confirming their inability to provide discriminative utility beyond a random guess on this balanced dataset. Similarly, CodeBERT and LineVul achieve slightly higher Accuracy (0.5163 and 0.5173) and perfect Recall (1.0000) but suffer from low Precision (0.5098 and 0.5103). This leads to middling F1-scores (0.6753 and 0.6757) and limited practical utility, as they cannot reliably discriminate between vulnerable and non-vulnerable functions, yielding a low MCC of 0.1212 and an AUC of 0.5144.
@@
More modern code-specific models show improved balance. UniXcoder and CodeT5 both outperform general text-embedding models and simple code-specific models, with UniXcoder reaching 0.5814 Accuracy and CodeT5 achieving 0.5854. These models strike a better trade-off between Recall and Precision, resulting in higher F1-scores (0.6822 and 0.6939, respectively) and improved discriminative power, as evidenced by their higher MCC values (0.2045 for UniXcoder and 0.2342 for CodeT5) and AUC scores (0.5796 and 0.5833, respectively). Most notably, CodeT5P delivers the strongest performance across all PLMs, achieving the highest Accuracy of 0.6037 along with a solid F1-score of 0.7075. It also reaches the highest MCC (0.2860) and AUC (0.6015) in the PLM category. Furthermore, the CodeT5P reaches FPR of 0.7496 and FNR of 0.0471, which implies that although CodeT5P can detect more vulnerable functions, it comes at the cost of wrongly detecting more clean functions. This indicates that CodeT5P, which incorporates identifier-aware training objectives and enriched representation learning, is particularly effective at capturing the multilingual semantic nuances required for function-level multilingual vulnerability detection. Overall, the results demonstrate that while traditional PLMs struggle with precision and overall Accuracy, task-aware models like CodeT5P offer significant improvements, making them more suitable for multilingual vulnerability detection scenarios at the function-level.